# School of Information and Network Security

### 03 Mar 2024 v2

@@ 4.2 Error Analysis
To gain a deeper understanding of the limitations of existing Large Language Models (LLMs) in handling complex tasks, we conducted a comprehensive error analysis on examples that performed poorly in the CPSDBench tasks. These errors can primarily be categorized into four types: triggering sensitive alerts, incorrect output formatting, failure to understand instructions, and erroneous content generation. These categories reflect the challenges LLMs face in comprehension, generation, and safety aspects. Figure 5 and figure 6 illustrate the distribution of error types in LLMs across two representative tasks: sentiment classification and telecommunications fraud detection.
@@ 4.2.2 Incorrect Output Format
Issues with output format represent one of the common error types for LLMs. In text classification and information extraction tasks, incorrect output formats directly impact the calculation of evaluation metrics. For example, the ChatGLM and ERNIE models incorporated additional explanatory information in their outputs, which did not comply with the task requirements for concise output formats. Moreover, when faced with open-ended questions, models tend to generate complete sentences rather than the required keywords or brief answers, leading to discrepancies between the predicted results and the standard answers.To assess the impact of neglecting formatting errors on the performance of LLMs) this study implemented an accuracy calculation method based on the principle of fuzzy matching. Through this method, we observed significant variations in the accuracy of multiple LLMs in tasks such as rumor detection and telecommunications fraud identification, with specific changes illustrated in figure  7 and figure  8. Notably, some models experienced an increase in accuracy by more than 20% after adopting fuzzy matching. This demonstrates that improving the formatting output capabilities of LLMs holds significant importance for tasks related to public safety.
Figure 7: After adopting fuzzy accuracy as the evaluation metric, the performance of LLMs in rumor detection tasks changed. The accuracy of the ChatGLM-4 model increased by 28%.
Figure 8: Following the implementation of fuzzy accuracy as an assessment criterion, the effectiveness of LLMs in telecommunications fraud detection tasks altered. The accuracy of the XVERSE2-13B model improved by 21%.
@@ 4.2.4 Errors in Output Content
Regarding errors in output content, especially in information extraction tasks, some models exhibited issues with missing key entities. Moreover, in text classification tasks, some models like Atom-1B tended to excessively output labels of a specific category, such as classifying instances as "normal" in telecommunication fraud detection tasks. This may reflect biases in the model’s training or overfitting to the training data distribution.Figures  9,  10, and  11 respectively illustrate the output bias of LLMs in three text classification tasks.
Figure 9: In the execution of sentiment classification tasks, the majority of models more frequently predict emotional tendencies as either "positive" or "negative," rather than opting for a "neutral" label.
Figure 10: Predictive bias of LLMs in rumor detection tasks
Figure 11: In telecommunications fraud detection tasks, LLMs exhibit a tendency to output "malicious" labels, which may lead to a higher false positive rate when applying them to this task.

### 21 Mar 2024 v3

@@ 4.2 Error Analysis
To gain a deeper understanding of the limitations of existing LLMs in handling complex tasks, we conducted a comprehensive error analysis on examples that performed poorly in the CPSDBench tasks. These errors can primarily be categorized into four types: triggering sensitive alerts, incorrect output formatting, failure to understand instructions, and erroneous content generation. These categories reflect the challenges LLMs face in comprehension, generation, and safety aspects. Figure 5 and figure 6 illustrate the distribution of error types in LLMs across two representative tasks: sentiment classification and telecommunications fraud detection.
@@ 4.2.2 Incorrect Output Format
Issues with output format represent one of the common error types for LLMs. In text classification and information extraction tasks, incorrect output formats directly impact the calculation of evaluation metrics. For example, the ChatGLM and ERNIE models incorporated additional explanatory information in their outputs, which did not comply with the task requirements for concise output formats. Moreover, when faced with open-ended questions, models tend to generate complete sentences rather than the required keywords or brief answers, leading to discrepancies between the predicted results and the standard answers. To assess the actual performance of LLMs after disregarding formatting errors, this study implemented an accuracy calculation method based on the principles of fuzzy matching. Through this method, we observed significant variations in the accuracy of multiple LLMs in text classification tasks, with specific changes presented in Figure 7. Notably, the accuracy of some models improved by more than 20% after adopting fuzzy matching for calculation. This demonstrates the significant importance of enhancing the formatting output capabilities of LLMs for tasks related to public security.
Figure 7: After adopting fuzzy accuracy as the evaluation metric, the performance of LLMs in text classification tasks changed. The accuracy improvement of models such as ChatGLM, Gemini, and XVERSE even exceeded 20%.
@@ 4.2.4 Errors in Output Content
Regarding errors in output content, especially in information extraction tasks, some models exhibited issues with missing key entities. Moreover, in text classification tasks, some models like Atom-1B tended to excessively output labels of a specific category, such as classifying instances as "normal" in telecommunication fraud detection tasks. This may reflect biases in the model’s training or overfitting to the training data distribution.Figures  8 illustrates the output bias of LLMs in three four classification tasks.
Figure 8: Predictive bias of LLMs in four text classification tasks. (1) In the execution of sentiment classification tasks, the majority of models more frequently predict emotional tendencies as either "positive" or "negative," rather than opting for a "neutral" label. (2) In telecommunications fraud detection tasks, LLMs exhibit a tendency to output "malicious" labels, which may lead to a higher false positive rate when applying them to this task.